Drift chamber HV system trouble debugging hints


  1. When the DCH cannot ramp up: Reboot the CAEN module in the EH (turn the key, wait a few seconds, turn it back).

  2. If it does not work: Diagnose which CAEN board is causing problems. To do this, open EPICS panel: DCH->High Voltage->More. This will show the four quadrants of the chamber, and the voltage in the various superlayers. Click "Sense", "Guard", or "Zipper" to select different wires. Look for values colored red or yellow--these are the ones that are in an error state. Click on the superlayer button (e.g., SL3_T), to open a small dialog box that will show which channel corresponds to that superlayer (e.g., PS_3:15). Now you need to figure out which board controls that channel. Use the following table to find this information: mapping. Now you should know which board is the problem.

  3. A quick fix that may work: Remove the 50 Ohm Lemo terminator from the offending board and swap it with the 50 Ohm terminator from another board. Lemo connectors are inherently flaky, so often a wiggle or a swap will make them work again.

  4. Another not-so-quick (~5 min.) fix to try: Reboot dch-mon (ramp down to 0). It could possibly improve the connection with the CAEN.

  5. If it still doesn't work, you need to look at the board: Remove the board , by following these steps.
    1. Turn off the CAEN with the key on the front of the rack.
    2. Go around to the back of the rack, and unplug the CAEN power cord from the 120V socket.
    3. Wait until the red LED (located on the back side of the CAEN crate, on the lower left beneath the cables) turns completely off. This may take several minutes.
    4. Once the LED is off, remove the power cable that connects the CAEN to the DCH HV. You need to unscrew the two screws and gently pull the cable out of the connection.
    5. Pull the board out of the rack.

  6. For the A944Ap boards, those in slots 2,3,4,5 (counting from 0): It's possible this board is still good, but that one of the chips came loose. So first try: Gently press all of the PROMs and driver chips into their sockets. These are at the back end of the board. You will likely hear a click as one of them snaps into place.

  7. You don't know if the board is bad or if the problem is somewhere downstream. So to isolate the board: Replace the CAEN's cable to DCH HV with a loop jump. You can find these in the DCH cabinet upstairs from the control room. It's a small cap that has two loops of wire. Attach one of these to the back of the CAEN. Now when you try to ramp up the DCH, if it fails you can be sure the problem is with the board, not farther downstream.

  8. Try the board again: Re-insert the board.
    1. Slide the board into the rack.
    2. Attach the cable, or loop jumper if you're debugging (be gentle, as the pins can easily bend).
    3. Make sure the Lemo connector is in place.
    4. Plug in the CAEN to the 120V socket.
    5. Turn the key counter-clockwise to Local On.

  9. If the original board still fails: Replace the old board with a spare. Note whether the board is an 833 or a 944. Find the appropriate spare from the IR2 Spares cabinet, which is on the back side of the DCH cabinet in the detector hall (key is in the Control Room close to the supply cabinet). Follow the steps in (8) to reinsert the board.

  10. Once you can ramp up successfully with the loop jumper, then you know the board works: Replace the jumper with the DCH HV cable.

  11. If you are using a new board, you need to: Reset the parameters for the new board. Open Epics to DCH->High Voltage->More R-Click->HV Configuration. Click on a board in the right column to view its set parameters.
    1. Compare all of the boards (4 sense, 2 guard, 1 zipper) to the reference plots. Since you've installed a new board, some of the settings will be wrong.
    2. Get superuser access. Go back to the main DCH Epics panel. The BaBar pilot must enable "Shifter Permission". Then you click on SU HV at work to get access.
    3. Update the settings. On the HV Configuration Epics panel, click "Update". In the GUI that opens, click on V0/V1 Update, and wait a minute for updates to install. Then click PS_1:I1 Update, and wait another minute.
    4. Now manually enter the necessary changes to settings (especially for boards 2, 3, and 4, the guard and zipper wires). From the HV Configuration panel, open up a board, and type the desired value to match the reference plots from part (a). Entering settings is a little tricky. Sometimes it works to enter the value in the cell beneath the given row. In other cases, you need to enter the value in the column to the right (with a "W" in the title) of the desired row--entering the value in the top cell usually works to propagate the settings all the way down the column.
    5. Once you have checked the settings on all the boards, sign off from superuser access, ramp up, and start getting data!
Last modified: Fri Mar 2 18:08:43 PST 2007